With respect to a particular question on a multiple choice test (having 4 alternatives with only 1 correct) a student  knows  and  therefore can elimlnate 3 of the 4 choices from consideration with probability 2/3 , can eliminate  2 of the 4 choices from consideration wid probability 1/6, can eliminate 1 choice from consideration wirh probability 1/9  and can eliminate none with probability 1/18 . If the student knows the answer , he answers correctly otherwise he guesses from among the choices not eliminated.


If the answer given by the student was found correct . then the probability that he knew the answer is a/b where 'a' and 'b' are relatively prime . find the value of (a+b).

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

a/b = (2/3)*1/(((2/3)"1)+0,5(1/6)+(1/9)(1/3)+(1/18)(1/4)) = 144/173, therefore a+b = 317